# Learning from Neighbors with PHIBP: Predicting Infectious Disease Dynamics in Data-Sparse Environments

### v1

@@ Abstract
Microbiome unseen species problems,
@@ 1 Introduction
Predicting disease prevalence in regions that have not yet reported cases—despite clear outbreaks in neighboring areas—is a problem most people now recognize from lived experience: when will it reach us, and how severe might it be? Turning this intuitive concern into reliable forecasts is a statistically challenging task. Historically, one can relate this to classes of count‑based models for both realized abundance (counts) and prediction of unseen numbers and types of species in the classic work of [2]. There Fisher, in answer to Corbett’s question about the expected number of distinct species of butterflies to be seen in a future sample, developed the logarithmic series distribution to describe the prevalence of individually observed species. This then translates into the remarkable work of [6], which empirically verifies the Negative Binomial distribution as a viable model for soil microorganism counts in the form of a mixed Poisson variable: the framework in [6] provides an empirical justification for a model in which the number of colonies is distributed according to a Poisson distribution, while the bacterial counts within each colony independently follow the logarithmic series distribution introduced by [2].
@@
Naturally, while many core elements of those works translate to modern datasets, they do not capture the notion of sharing information across groups, nor matters of count sparsity. In particular, while the prevalence of zeros is common in modern datasets, we further highlight the notion of sparse co-occurrence, where samples across groups do not share many common species. These challenges, along with modeling complex multivariate count distributions, are all addressed by the Poisson Hierarchical Indian Buffet Process (PHIBP) [5] as applied to complex microbiome sampling models. Here, we focus on its ability to pool information across groups to handle zero observations, leading to credible predictive and inferential modeling for infectious disease prediction. As the theory of PHIBP has been detailed previously, our focus is on methodology and computational inference within the present context of disease prediction.
@@ 2 The PHIBP model
where $Y_{l}$ indexes a global catalogue of disease types and $N^{(i)}_{j,l}\in\{0,1,2,\dots\}$ is the count of disease $Y_{l}$ in sample $i$ from region $j$. The PHIBP specifies a hierarchical prior on the latent mean abundance rates that generate these counts. At the top level, we draw a global completely random measure
@@
where $\tau_{0}$ is a Lévy measure on $(0,\infty)$ and $F_{0}$ a non-atomic probability measure on a Polish space $\Omega$ of disease labels. The jumps $(\lambda_{l})$ are interpreted as global mean rates of the diseases $(Y_{l})$ across all regions.
@@
with Lévy density $\tau_{j}$ on $(0,\infty)$. Conditionally on $B_{0}=\sum_{l}\lambda_{l}\delta_{Y_{l}}$ we may write
@@
where $(\sigma_{j}(t):t\geq 0)$ is a subordinator with Lévy density $\tau_{j},$ satisfying for $s<t,$ $\sigma_{j}(t)-\sigma_{j}(s)\overset{d}{=}\sigma_{j}(t-s)$ independent of $\sigma_{j}(s)$. Thus $\lambda_{l}$ is the global rate of disease $Y_{l}$, and $\sigma_{j,l}(\lambda_{l})$ is its local mean rate in region $j$.
Throughout, as in [5], we denote by $\psi_{j}(\cdot)$ and $\Psi_{0}(\cdot)$ the Laplace exponents associated with the Lévy measures $\tau_{j}$ and $\tau_{0}$, respectively. In particular, for each $j$ and exposure $\gamma_{j}>0$, $\mathbb{E}[e^{-\gamma_{j}\sigma_{j}(\lambda)}]=e^{-\lambda\psi_{j}(\gamma_{j})}$, where
$\psi_{j}(\gamma_{j})\;=\;\int_{(0,\infty)}\big(1-e^{-\gamma_{j}s}\big)\,\tau_{j}(ds),$
and similarly
@@
where we assume that $\int_{0}^{\infty}\min(s,1)\tau_{j}(s)ds<\infty$ ensuring the finiteness of $B_{j}(\Omega)$ for $j=0,\ldots,J.$ We will only use these via evaluations such as $\psi_{j}\!\big(\sum_{i=1}^{M_{j}}\gamma_{i,j}\big)$ and $\Psi_{0}\big(\sum_{j=1}^{J}\psi_{j}(\cdot)\big)$, exactly as in the PHIBP construction of [5]. Furthermore setting $\psi^{(c)}_{j}(\gamma_{j})=\int_{0}^{\infty}s^{c}e^{-\gamma_{j}s}\tau_{j}(s)ds$ we have the mixed truncated Poisson distributions with law denoted as $\mathrm{MtP}(\tau_{j},\gamma_{j})$ and probability mass function for $c=1,2,\ldots,$ given by
$\frac{\gamma^{c}_{j}\psi^{(c)}_{j}(\gamma_{j})}{\psi_{j}(\gamma_{j})c_{!}}$
The PHIBP $(Z^{(i)}_{j},i\in[M_{j},j\in[J])$ is specified as a mixed Poisson processes driven by the local CRMs. For each $j$ and $i\in[M_{j}]$,
@@
This is the basic mixed Poisson representation: each observed count is Poisson with a random intensity given by the exposure $\gamma_{i,j}$ times the local rate $\sigma_{j,l}(\lambda_{l})$.
@@
which governs the rate of species (atoms) observed in region $j$ after aggregating over its $M_{j}$ samples. Writing $N_{j,l}:=\sum_{i=1}^{M_{j}}N^{(i)}_{j,l}$ for the total count of disease $Y_{l}$ in region $j$, we have
@@
$\Big(\sum_{i=1}^{M_{j}}Z^{(i)}_{j},\;j\in[J]\Big)\;\overset{d}{=}\;\Big(\sum_{\ell=1}^{\varphi}\tilde{N}_{j,\ell}\,\delta_{\tilde{Y}_{\ell}},\;j\in[J]\Big)\;\overset{d}{:=}\;\Big(\sum_{\ell=1}^{\varphi}\big[\sum_{k=1}^{X_{j,\ell}}C_{j,k,\ell}\big]\delta_{\tilde{Y}_{\ell}},\;j\in[J]\Big).$ (2.3)
admits an exact compound Poisson representation in terms of a random number of latent (sub)-species level clusters and their counts. Here $\varphi$ is the random number of distinct species that appear across all regions, and has a Poisson distribution with mean $\Psi_{0}\big(\sum_{j=1}^{J}\psi_{j}(\sum_{i=1}^{M_{j}}\gamma_{i,j})\big)$. For each observed species $\ell$, there is a global posterior mean rate $H_{\ell}>0$ and a corresponding total number of OTU-level clusters $X_{\ell}:=\sum_{j=1}^{J}X_{j,\ell}$; jointly, $(H_{\ell},X_{\ell})$ has the mixed Poisson–MtP structure of [5], with $X_{\ell}\sim\mathrm{MtP}\big(\tau_{0},\sum_{j=1}^{J}\psi_{j}(\sum_{i=1}^{M_{j}}\gamma_{i,j})\big)$ and $H_{\ell}$ conditionally distributed with density proportional to $\lambda^{X_{\ell}}\exp\!\{-\lambda\sum_{j=1}^{J}\psi_{j}(\sum_{i=1}^{M_{j}}\gamma_{i,j})\}\tau_{0}(\lambda)$. Given $X_{\ell}=x_{\ell}$, the allocation of OTU clusters across regions is multinomial, $(X_{1,\ell},\dots,X_{J,\ell})\mid X_{\ell}=x_{\ell}\sim\mathrm{Multinomial}(x_{\ell};q_{1},\dots,q_{J})$, with $q_{j}\propto\psi_{j}\big(\sum_{i=1}^{M_{j}}\gamma_{i,j}\big)$. Finally, given these allocations and the local Lévy measures $\tau_{j}$, the OTU-level cluster sizes $C_{j,k,\ell}$ are independent mixed truncated Poisson variables $C_{j,k,\ell}\sim\mathrm{MtP}\big(\tau_{j},\sum_{i=1}^{M_{j}}\gamma_{i,j}\big)$, so that each regional total $\tilde{N}_{j,\ell}=\sum_{k=1}^{X_{j,\ell}}C_{j,k,\ell}$ is a sum of i.i.d. MtP components driven by the global rate $H_{\ell}$.

### v2

@@ Abstract
Unseen species prediction,
@@ 1 Introduction
Predicting disease prevalence in regions that have not yet reported cases—despite clear outbreaks in neighboring areas—is a problem most people now recognize from lived experience: when will it reach us, and how severe might it be? Turning this intuitive concern into reliable forecasts is a statistically challenging task. Historically, one can relate this to classes of count-based models for both realized abundance (counts) and prediction of unseen numbers and types of species in the classic work of [3]. There Fisher, in answer to Corbett’s question about the expected number of distinct species of butterflies to be seen in a future sample, developed the logarithmic series distribution to describe the prevalence of individually observed species. This then translates into the remarkable work of [7], which empirically verifies the Negative Binomial distribution as a viable model for soil microorganism counts in the form of a mixed Poisson variable: the framework in [7] provides an empirical justification for a model in which the number of colonies is distributed according to a Poisson distribution, while the bacterial counts within each colony independently follow the logarithmic series distribution introduced by [3].
@@
Naturally, while many core elements of those works translate to modern datasets, they do not capture the notion of sharing information across groups, nor matters of count sparsity. In particular, while the prevalence of zeros is common in modern datasets, we further highlight the notion of sparse co-occurrence, where samples across groups do not share many common species. These challenges, along with modeling complex multivariate count distributions, are all addressed by the Poisson Hierarchical Indian Buffet Process (PHIBP) [6] as applied to complex microbiome sampling models. Here, we focus on its ability to pool information across groups to handle zero observations, leading to credible predictive and inferential modeling for infectious disease prediction.
The present work is part of a three-paper arc. The foundational work [6] establishes PHIBP as a comprehensive statistical framework for hierarchical count data. Its theoretical complement [5] resolves long-standing problems in coagulation–fragmentation duality, revealing deep structural properties of the framework that extend to continuous-time processes and complex ancestral constructions. This application demonstrates PHIBP’s practical capabilities for disease surveillance while simultaneously showcasing the broader extensibility of the trilogy: the same underlying architecture accommodates both the esoteric theoretical developments and domain-specific customizations amenable to future AI/LLM frameworks for intelligent model specification.
Before detailing the model, we first outline several architectural innovations that distinguish PHIBP and motivate its use for disease prediction. As the theory of PHIBP has been detailed previously, our focus is on methodology and computational inference within the present context of disease prediction.
2 Architectural Innovations of PHIBP
2.1 Absolute abundance vs. compositional rates
A fundamental design choice in count-based modeling is whether to work with absolute abundances (raw counts or rates) or relative abundances (proportions summing to unity within each sample). While compositional approaches—dominant in modern microbiome and ecological analyses—offer certain computational conveniences, they suffer from the closure problem: because relative abundances must sum to one, all components are mechanically constrained such that an increase in one component necessitates a proportional decrease in others, even when their absolute counts remain unchanged. This induces spurious negative correlations among components that reflect the mathematical constraint rather than genuine biological or epidemiological relationships. For infectious disease surveillance, where distinguishing true co-occurrence patterns from measurement artifacts is critical, this confounding is particularly problematic.
Consider a concrete example: two counties each report 100 total disease cases in Year 1, with County A showing 60 cases of Disease X and 40 of Disease Y, while County B shows 50 cases each. In Year 2, both counties experience identical 20-case increases in Disease X (reaching 80 and 70 cases respectively), while Disease Y remains stable at 40 and 50 cases. The absolute data reveal a shared trend: Disease X is spreading identically in both locations. Yet compositional methods see County A shifting from $60\%$$40\%$ to $67\%$$33\%$ and County B from $\%/5050\%$ to $58\%$$42\%$—artificially suggesting that Disease Y is declining when its absolute burden is stable. When data contain many zeros, the distortions compound: a single new case drastically alters all proportions, and counties with zero cases cannot be meaningfully compared under relative-abundance frameworks.
The absolute-abundance perspective has deep roots in classical statistical ecology. Fisher, Corbet, and Williams (1943) modeled butterfly species abundances using the logarithmic series distribution for the frequency of species with $1,2,3,\dots$ individuals in a fixed sample, estimating the underlying abundance parameter from the observed counts rather than working with normalized compositional proportions. This foundational work then translates into the remarkable empirical investigations of [7], which verified the negative binomial as a viable model for soil microorganism counts in the form of a mixed Poisson variable. The framework in [7] provides an empirical justification for a model in which the number of colonies is distributed according to a Poisson distribution, while the bacterial counts within each colony independently follow the logarithmic series distribution introduced by Fisher. Both approaches model absolute counts and treat each species’ abundance as an independent realization, avoiding the sum-to-one constraint. This classical tradition recognizes that the proper modeling target is the intensity (mean count per unit exposure), not the proportion of a fixed total. The PHIBP foundations are strongly connected to, and influenced by, Pitman’s manuscript [10], which revisits this Fisher–McCloskey perspective and highlights it as a neglected strand in the species sampling literature.
The PHIBP framework inherits and extends this absolute-abundance paradigm to hierarchical settings. At the global level, each disease type $\ell$ is assigned an absolute rate $\lambda_{\ell}>0$ representing its mean intensity. Locally, county $j$ experiences a transformed rate $\sigma_{j,\ell}(\lambda_{\ell})$ obtained via a subordinator process that modulates the global rate according to region-specific characteristics captured by the Lévy measure $\tau_{j}$. Crucially, these rates are not constrained to sum to any fixed value: a high rate for Disease A in County 1 does not mechanically reduce the inferred rate for Disease B, nor does it affect Disease A’s rate in County 2. The subordinator architecture preserves marginal independence of disease intensities [6].
This absolute-abundance framework has immediate implications for interpreting zero counts. Under compositional approaches, a county reporting zero cases of Disease A and five of Disease B is forced into the representation (0%, 100%), which is undefined when computing log-ratios and creates computational difficulties. More fundamentally, compositional methods cannot distinguish sampling zeros—where Disease A is present at low rate but unobserved—from structural zeros, where it genuinely cannot occur. PHIBP’s absolute-rate parameterization accommodates both: a county can have small but nonzero $\sigma_{j,\ell}(\lambda_{\ell})$ yielding zero observed counts via Poisson variability, or $\sigma_{j,\ell}(\lambda_{\ell})\approx 0$ indicating true absence. The posterior distribution over these rates, informed by hierarchical borrowing, assigns appropriate uncertainty and enables probabilistic prediction of emergence in currently zero-count locations.
By returning to Fisher’s absolute-abundance tradition and extending it through subordinator-driven hierarchical rates, PHIBP aligns the mathematical structure with the epidemiological question: not “what proportion of cases are Disease X?” but “how many cases of Disease X do we expect per unit population?”
2.2 Thinning vs. exchangeable sampling
A second architectural choice is whether to generate observations by thinning a Poisson random measure or by exchangeable sampling from a normalized random measure. In classical exchangeable partition models, one first constructs a random probability measure GG and then obtains data by drawing a fixed sample of size $n$ i.i.d. from $G$. This induces elegant partition laws, but it implicitly treats $n$ as the fundamental quantity: the model is conditioned on the total number of individuals or reads observed. For sequencing-based surveillance or routine reporting systems, this conditioning is artificial. The total count is itself a random outcome of the measurement process, driven by laboratory throughput, reporting practices, or population size, rather than a fixed design parameter.
The PHIBP framework instead follows the thinning perspective inherent in its mixed Poisson construction. Starting from global mean intensities on the disease types, it generates local group-level intensities, and only then produces observed counts by Poisson sampling, with rates scaled by the relevant exposure (e.g. population-at-risk, test volume, or reporting effort). In this view, the primary random objects are the underlying intensities and the exposure; the total count is a derived random quantity. This aligns more naturally with infectious disease settings, where one does not control the number of cases, but rather observes whatever the system produces over a given period.
This distinction has several practical consequences. First, thinning provides a transparent representation of sampling zeros versus structural zeros: a disease may have strictly positive but small intensity in a county and still yield zero observed cases by chance, whereas an intensity that is shrunk toward zero across the hierarchy corresponds to genuine absence. Exchangeable sampling on normalized measures, by contrast, operates on proportions and a fixed $n$, and has no native mechanism to represent the random occurrence of zeros at fixed exposure. Second, thinning leads directly to additive prediction rules. Adding a new surveillance period or a new region corresponds to drawing new Poisson counts at the same latent intensities; the three-way decomposition into (i) completely new disease types, (ii) newly observed manifestations of existing types, and (iii) additional counts for already-observed types follows from standard Poisson calculus, as in the PHIBP prediction rules.
Finally, thinning is what makes the PHIBP-based architecture computationally usable in our setting: the mixed Poisson and compound Poisson representations depend critically on viewing the data as Poisson thinnings of underlying rate processes, rather than as exchangeable samples of fixed size. This is precisely what allows us to derive explicit posterior and predictive laws for sparse, hierarchically structured counts.
Taken together, these choices place PHIBP firmly in the absolute-abundance tradition of Fisher, Corbet, and Quenouille, where intensities and mixed Poisson structure are the primary modelling objects, while at the same time marking a genuine paradigm shift. The framework extends that tradition to hierarchical, multi-group settings with explicit rate coupling and exact joint laws for complex dependence structures—features that are not available in classical relative-abundance or purely exchangeable formulations.
Remark 2.1.
For a deeper look at thinning-based sampling mechanisms as primary, exchangeability as a subsequent design choice, and the flexibility to design hierarchical models on abstract Polish spaces using $h$-biased measures with domain-specific interpretations, readers are invited to consult [5, Section 7].
2.3 Extensibility of the PHIBP framework
Although not emphasized as a central structural feature, the PHIBP framework subsumes the capabilities of hierarchical compositional models such as the Hierarchical Dirichlet Process (HDP) [13] and its extensions. In other words, PHIBP can accomplish everything these classical models can do—but the converse is not true. This generality is made strikingly apparent in the coagulation–fragmentation duality work [5, see e.g. Sections 1.3, 3, and Proposition 2.5].
The present exposition should be viewed as part of a three-paper arc. The foundational work [6] establishes PHIBP as a comprehensive statistical framework for species sampling models and hierarchical count data. Its theoretical complement [5] resolves long-standing problems in coagulation–fragmentation duality for both static and continuous-time settings, operating primarily within probability theory while revealing substantial capabilities for complex applications. The current disease surveillance application demonstrates aspects of this framework in action, though it operates within the essential skeleton of PHIBP itself. Our intent—as in the foundational work—is to provide a principled building block on which domain experts can construct bespoke extensions tailored to their specific problems.
Section 7 of [5] is particularly relevant here: it describes how to design customized PHIBP variants by directly specifying the behavior of Poisson random measure components in ways that carry concrete meaning for epidemiologists, ecologists, or other domain practitioners—rather than selecting a modeling component, such as a Gamma process or stable subordinator, merely because it yields a tractable closed-form distribution. As developed in that work, PHIBP variants now manifest as a $4\times J$ component dual system—simultaneously encoding fine and coarse partitions, and fragmentation and coagulation operators—that naturally elevates to continuous-time models and supports sophisticated ancestral constructions such as structured Ancestral Recombination Graph (ARG) and Time to the Most Recent Common Ancestor (TMRCA) frameworks. Within this cloud-duality setup, domain experts may encode substantive structure directly through the choice of state spaces $S_{j}$, mean measures, and $h_{j}$-biasing functions on point clouds, yielding models whose components are interpretable in scientific terms rather than imposed for distributional convenience. With this architecture, modeling choices can be driven by interpretability and domain insight rather than mathematical tractability alone, and the framework is not bound to standard sampling schemes such as exchangeability [6]. Notably, these capabilities align naturally with the evolving potential of AI/LLM systems for model specification, allowing domain experts to translate conceptual knowledge—for example, about ARG topologies, recombination counts, or TMRCA behavior—into explicit prior structure within a principled probabilistic framework.
@@ 3 The PHIBP model
Modeling objectives and contribution. This chapter applies the PHIBP framework to infectious disease prediction, demonstrating its computational implementation in a setting where many county-disease pairs exhibit zero observed counts. The primary inferential targets are the latent disease-specific rates—global rates $H_{\ell}$ and local rates $\sigma_{j,\ell}(\lambda_{\ell})$ for each region $j$ and disease type $\ell$—which govern the intensity of disease occurrence. From these rates, we derive posterior distributions for alpha- and beta-diversity measures and construct predictive distributions for diseases not yet observed in specific counties. The hierarchical structure enables information borrowing: counties with zero cases of a disease can still receive informative posteriors by pooling evidence from neighboring regions and the global disease catalog. While the foundational theory is developed in [6], the present application demonstrates that the PHIBP framework operates without modification on epidemiological data that appear structurally distinct from the microbiome context, revealing the framework’s generality and flexibility. Furthermore, the architecture admits substantial extensions—hierarchical geographic structures, dynamic surveillance, and integration with domain-specific covariates—that we outline in Section 5, positioning PHIBP as a versatile building block for hierarchical count modeling across scientific domains.
We introduce the PHIBP framework for modeling hierarchical count data. For a comprehensive treatment of the underlying technical foundations—including Poisson random measures, completely random measures, and Lévy subordinators—the reader may consult standard references such as [1, 12]. Furthermore, while we aim for a self-contained exposition, the reader is encouraged to consult the more comprehensive treatment in the primary PHIBP reference [6].
Before detailing the hierarchical structure, we briefly recall the foundational object. A Poisson random measure (PRM) $N$ over a Polish space $\Omega$ is a random measure with the following properties:
  1. (i)
    For disjoint measurable sets $A,B\subset\Omega$, the random variables $N(A)$ and $N(B)$ are independent.
  2. (ii)
    For any measurable set $A$, $N(A)\sim\mathrm{Poisson}(\nu(A))$, where $\nu$ is a $\sigma$-finite measure on $\Omega$ called the mean measure, satisfying $\mathbb{E}[N(A)]=\nu(A)$.
While random measures may seem daunting, the key insight is that they can be thought of in terms of random variables or random vectors: $N(A)$ is simply a Poisson random variable whose intensity depends on the set $A$. This perspective makes them interpretable and easily incorporated into various applications. This same interpretability carries over to functionals of PRMs, such as completely random measures (CRMs), which we describe next. Intuitively, a CRM is a random intensity measure constructed by placing random jumps at random locations on a Polish space, with the jumps governed by a Lévy measure. The marginal distributions of such measures are based on familiar infinitely divisible random variables such as Gamma, stable, and generalized Gamma, as well as more exotic choices. A natural way of characterizing such variables is via Laplace transforms, and this together with independence over disjoint sets is enough to characterize these processes.
The ability to operate on abstract Polish spaces is not only for theoretical flexibility but enables bespoke designs for applications by domain experts, as illustrated in Section 7 of [5]. Here, we use $\Omega$ generically throughout. For disease applications, $\Omega$ is a finite or countable catalogue of circulating disease types (pathogens, strains, etc.) in the population; more abstractly, it indexes data labels or types that inherit structure from the hierarchical specification.
A completely random measure (CRM) is constructed from a PRM on the product space $(0,\infty)\times\Omega$ by integrating out the jump component. Specifically, let $N$ be a PRM on $(0,\infty)\times\Omega$ with mean measure $\nu(ds,dy)=\tau(ds)\,F_{0}(dy)$, where $\tau$ is a Lévy measure on $(0,\infty)$ satisfying $\int_{(0,\infty)}\min(s,1)\,\tau(ds)<\infty$, and $F_{0}$ is a $\sigma$-finite measure on $\Omega$. Throughout, we write Lévy measures in density form, for example $\tau(ds)=\tau(s)\,ds$, and adopt this compact notation for convenience.
Then the CRM is the random atomic measure
$B_{0}\;=\;\int_{(0,\infty)\times\Omega}\lambda\,N(d\lambda,dy)\;=\;\sum_{l\geq 1}\lambda_{l}\,\delta_{Y_{l}},$
where the sum ranges over the (almost-surely countable) atoms $(\lambda_{l},Y_{l})$ of $N$, and we write $B_{0}\sim\mathrm{CRM}(\tau,F_{0})$. The Lévy measure $\tau$ governs the distribution of jump sizes $(\lambda_{\ell})$, and its Laplace exponent is $\psi(\gamma):=\int_{(0,\infty)}(1-e^{-\gamma s})\,\tau(ds)$ for $\gamma\geq 0$.
For the PHIBP framework developed here, we choose $F_{0}$ to be a probability measure on $\Omega$—this means we do not need to track a normalizing constant in the mean intensity.
@@
where $Y_{l}$ indexes a global catalogue of disease types and $N^{(i)}_{j,l}\in\{0,1,2,\dots\}$ is the observed count of disease type $Y_{l}$ in sample $i$ from region $j$. The PHIBP specifies a hierarchical prior on the latent mean abundance rates that generate these counts. At the top level, we draw a global completely random measure
@@
where $\tau_{0}$ is a Lévy measure on $(0,\infty)$ and $F_{0}$ a non-atomic probability measure on the Polish space $\Omega$ of disease types. The jumps $(\lambda_{l})$ are interpreted as global mean rates of the diseases $(Y_{l})$ across all regions.
@@
with Lévy measure $\tau_{j}$ on $(0,\infty)$. When $B_{0}$ is not a probability measure, this is well-defined as follows: we condition on the realized atoms $(\lambda_{\ell},Y_{\ell})$ of $B_{0}$ and construct the local jumps accordingly. Conditionally on $B_{0}=\sum_{l}\lambda_{l}\delta_{Y_{l}}$ we may write
@@
where $\sigma_{j,l}(\lambda_{l})=\sum_{k\geq 1}s_{j,k,l}$ and the atoms $(s_{j,k,l})_{k\geq 1}$ are the jumps of the subordinator $\sigma_{j}$ restricted to the interval of length $\lambda_{l}$. More formally, $(\sigma_{j}(t):t\geq 0)$ is a subordinator with Lévy measure $\tau_{j}$, satisfying for $s<t$, $\sigma_{j}(t)-\sigma_{j}(s)\overset{d}{=}\sigma_{j}(t-s)$ independent of $\sigma_{j}(s)$. Thus $\lambda_{l}$ is the global rate of disease type $Y_{l}$, and $\sigma_{j,l}(\lambda_{l})=\sum_{k\geq 1}s_{j,k,l}$ is its local mean rate in region $j$, decomposed into finer-grained subspecies-level intensities $(s_{j,k,l})$. This construction—using $B_{0}$ as the directing measure for $B_{j}$—is what allows the local rates to inherit the disease labels $(Y_{l})$ from the global level while modulating their intensities via the subordinator, with the additional structure that each disease type $Y_{l}$ harbors internal heterogeneity captured by the subspecies markers indexed by $k$.
Throughout, as in [6], we denote by $\psi_{j}(\cdot)$ and $\Psi_{0}(\cdot)$ the Laplace exponents [1] associated with the Lévy measures $\tau_{j}$ and $\tau_{0}$, respectively. For a subordinator $\sigma_{j}$ with Lévy measure $\tau_{j}$, the Laplace exponent $\psi_{j}$ is defined by
$\mathbb{E}[e^{-\gamma_{j}\sigma_{j}(t)}]=e^{-t\psi_{j}(\gamma_{j})},$
where
$\psi_{j}(\gamma_{j})\;=\;\int_{(0,\infty)}\big(1-e^{-\gamma_{j}s}\big)\,\tau_{j}(ds).$
Similarly, for the global subordinator $\sigma_{0}$ with Lévy measure $\tau_{0}$,
@@
The Laplace exponent $\psi_{j}(\gamma_{j})$ depends only on $\gamma_{j}$ and the Lévy measure $\tau_{j}$, not on any particular time $t$ or magnitude $\lambda$. We assume that $\int_{0}^{\infty}\min(s,1)\tau_{j}(ds)<\infty$ ensuring the finiteness of $B_{j}(\Omega)$ for $j=0,\ldots,J$. In practice, we evaluate $\psi_{j}$ at aggregated exposures such as $\psi_{j}\!\big(\sum_{i=1}^{M_{j}}\gamma_{i,j}\big)$ and compose exponents as $\Psi_{0}\big(\sum_{j=1}^{J}\psi_{j}(\cdot)\big)$, exactly as in the PHIBP construction of [6].
Furthermore, setting
$\psi^{(c)}_{j}(\gamma_{j})=\int_{0}^{\infty}s^{c}e^{-\gamma_{j}s}\,\tau_{j}(ds),$
we have the mixed truncated Poisson distributions with law denoted by $\mathrm{MtP}(\tau_{j},\gamma_{j})$ and probability mass function, for $c=1,2,\ldots$, given by
$\frac{\gamma^{c}_{j}\,\psi^{(c)}_{j}(\gamma_{j})}{\psi_{j}(\gamma_{j})\,c!},$
where
$\sum_{c=1}^{\infty}\frac{\gamma^{c}_{j}\,\psi^{(c)}_{j}(\gamma_{j})}{c!}=\psi_{j}(\gamma_{j}).$
The PHIBP $(Z^{(i)}_{j},i\in[M_{j}],j\in[J])$ is specified as a mixed Poisson point process (PoiP) driven by the local CRMs; a Poisson point process is the random counting measure generated by a CRM, counting events according to the intensities encoded in the measure. For each $j$ and $i\in[M_{j}]$,
@@
This is the basic mixed Poisson representation: each observed count $N^{(i)}_{j,l}$ is an outcome of a Poisson random variable with mean $\gamma_{i,j}\sigma_{j,l}(\lambda_{l})$, the product of exposure and local intensity.
@@
which governs the rate of distinct disease types (atoms) observed in region $j$ after aggregating over its $M_{j}$ samples. Writing $N_{j,l}:=\sum_{i=1}^{M_{j}}N^{(i)}_{j,l}$ for the total count of disease type $Y_{l}$ in region $j$, we have
@@
$\Big(\sum_{i=1}^{M_{j}}Z^{(i)}_{j},\;j\in[J]\Big)\;\overset{d}{=}\;\Big(\sum_{\ell=1}^{\varphi}\tilde{N}_{j,\ell}\,\delta_{\tilde{Y}_{\ell}},\;j\in[J]\Big)\;\overset{d}{:=}\;\Big(\sum_{\ell=1}^{\varphi}\big[\sum_{k=1}^{X_{j,\ell}}C_{j,k,\ell}\big]\delta_{\tilde{Y}_{\ell}},\;j\in[J]\Big)$ (3.3)
admits an exact compound Poisson representation in terms of a random number of latent (sub)-species level clusters and their counts. Here $\varphi$ is the random number of distinct disease types that appear across all regions, and has a Poisson distribution with mean
$\Psi_{0}\big(\sum_{j=1}^{J}\psi_{j}(\sum_{i=1}^{M_{j}}\gamma_{i,j})\big).$
For each observed disease type $\ell$, there is a global posterior mean rate $H_{\ell}>0$. The number of operational taxonomic units (OTUs)—representing the finest-grained hierarchical classification within each disease type—in region $j$ is denoted $X_{j,\ell}$, with total across all regions $X_{\ell}:=\sum_{j=1}^{J}X_{j,\ell}$. Jointly, $(H_{\ell},X_{\ell})$ has the mixed Poisson–MtP structure of [6], with $X_{\ell}\sim\mathrm{MtP}\big(\tau_{0},\sum_{j=1}^{J}\psi_{j}(\sum_{i=1}^{M_{j}}\gamma_{i,j})\big)$ and $H_{\ell}\mid X_{\ell}=x_{\ell}$ conditionally distributed with density proportional to
$\lambda^{x_{\ell}}\exp\!\{-\lambda\sum_{j=1}^{J}\psi_{j}(\sum_{i=1}^{M_{j}}\gamma_{i,j})\}\tau_{0}(\lambda).$
Given $X_{\ell}=x_{\ell}$, the allocation of OTU clusters across regions is multinomial,
$(X_{1,\ell},\dots,X_{J,\ell})\mid X_{\ell}=x_{\ell}\sim\mathrm{Multinomial}(x_{\ell};q_{1},\dots,q_{J}),$
with $q_{j}\propto\psi_{j}\big(\sum_{i=1}^{M_{j}}\gamma_{i,j}\big)$. Finally, given these allocations and the local Lévy measures $\tau_{j}$, the OTU-level cluster sizes $C_{j,k,\ell}$ are independent mixed truncated Poisson variables
$C_{j,k,\ell}\sim\mathrm{MtP}\big(\tau_{j},\sum_{i=1}^{M_{j}}\gamma_{i,j}\big),$
so that each regional total $\tilde{N}_{j,\ell}=\sum_{k=1}^{X_{j,\ell}}C_{j,k,\ell}$ is a sum of i.i.d. MtP components driven by the global rate $H_{\ell}$. The joint vector of random counts $(\tilde{N}_{j,\ell},j\in[J])$ is independent over $\ell\in[\varphi]$ and is shown to have an explicit, albeit more complex, form within the MtP family in [6, Proposition 4.3], and plays important roles in [5, Section 2].
Remark 1.
The jump decomposition $(s_{j,k,\ell})$ underlying the local rates $\sigma_{j,\ell}(\lambda_{\ell})$ admits representation in terms of $h$-biased measures, a framework developed in Section 7 of [5]. Specifically, one may write $s_{j,k,\ell}=h_{j}(\mathbf{e}_{j,k,\ell})$, where $(\mathbf{e}_{j,k,\ell})$ are points of a Poisson random measure on more general spaces $S_{j}$, and $h_{j}$ are non-negative functions. This flexibility allows domain experts to encode substantive structure directly—through the choice of $S_{j}$ and $h_{j}$—rather than selecting components solely for mathematical convenience. In disease surveillance, for instance, $S_{j}$ might index transmission pathways, and $h_{j}$ might encode geographic proximity or demographic factors, yielding interpretable hierarchical structures tailored to epidemiological questions.